OpenStudy (rianiscool):

need give asap will give a medal need help asap will give a medal Jorge's rectangular garden was 17 feet long and had a perimeter of 54 feet. He made his garden 5 feet longer. What is the area of his new garden? A. 22 square feet B. 64 square feet C. 85 square feet D. 220 square feet

OpenStudy (rianiscool):

is it D i think

OpenStudy (rianiscool):

@iGreen

OpenStudy (gorv):

perimeter= 2*length+2*width

OpenStudy (gorv):

54=2*17+2*width 54=34+2*width 54-34=2*width 20=2*width width=20/2=????

OpenStudy (rianiscool):

so is it D??

OpenStudy (gorv):

now new length=17+5

OpenStudy (gorv):

yeah

OpenStudy (rianiscool):

ok thanks

OpenStudy (rianiscool):

gave you a medal to @gorv

OpenStudy (gorv):

thnks :)